#12

Chapter 3 Problem 3.05 12 of 16> V is a vector 24.8 units in magnitude and points at an angle of 234 above the negative z axis No elements selected IM
s a vector 24.8 units in magnitude and points at an angle of 4 above the negative z axis Request Answer Part C Use V, and V to obtain (again) the direction ot V URL 6 7 3
Answer #1

个Vy Vx

This is how vector V will look like. Te angle which V vector makes with negative x axis is 23.4 degree.

NOTE - The image is not very clear. I see the magnitude as 24.8 units

Part B

Vx = - 24.8 * cos 23.4 = -22.76 units

Vy = 24.8 * sin 23.4 = 9.849 units

-----------------------------------------------------------

so,

Part C

V = sqrt (22.762 + 9.8492)

V = 24.799 units

----------------------------------------------------------

Part D

direction of V

theta = arctan (9.849/22.76)

theta = 23.4 degree